Calculate the pH for each of the following cases in the titration of 25.0 mL of 0.190 M pyridine, C5H5N(aq) with 0.190 M HBr(aq):

a.) Before addition of HBr?

b.) After addition of 12.5 mL of HBr?

c.) After addition of 15.0 mL of HBr?

d.) After addition of 25.0 mL of HBr?

e.) After addition of 34.0 mL of Hbr?
Answer #1

pH formulas for salts weak acid + strong baise salt tho pH at { pk w tp ka +Log e] strong acid + weak base salt tho pH = [I < Ky & SHANE 17X10 9 altsN+ HO H H + oH 0.19 A O -* 0.19 * [est Nit] Lott] [5H₃N] +X 1-2810a 0-19- 1-28109 (0.19 -x) = x²® stan molarity 2 0.19M volume 225 ml=0.0252 HBr molarity a o.19M volume 12.5ml=0.01252 no of moles o.19 X 0.0 125 -0.002375HBr molarity = o- LAM GHN Molarity oLaM volume = 15ml = 0.0152 volume - 25 ml=0.0252 no- of moles= 0.19 x 0.015 no ob moles=H₂N Molarity - O-lam po HBr molarity 0.19M volume = 25ml = 0.0252 volume = 25ml=0.0252 no ob moleso. L9X 0.025 no- of moles-0GHSN Molarity o-LaM HBr Molarity = 0.19M volume = 25m1 = 0.0252 volume = 34m120.0342 no of molesa 8. L9X0.025 no of moles - 0